OpenStudy (anonymous):

Heres a story problem i need help solving. Two cars travel at right angles to each other from an intersection until they are 39 miles apart. At this point, one car has gone 29 miles farther than the other. How far did the slower car travel?

OpenStudy (amistre64):

its a right triangle

OpenStudy (amistre64):

hyp= 39 leg = 29 other leg = ?

OpenStudy (amistre64):

i gotta learn to read the problems all the way thru ... leg = 29+x other leg = x

OpenStudy (amistre64):

i fyou recall the pythag thrm; this should be a breeze

OpenStudy (anonymous):

39^ = 29^2 + x^2 Solve for x.

OpenStudy (amistre64):

you did the same as; but its 29 further than x; so (x+29)^2

OpenStudy (anonymous):

YOu are right, I did do the same thing, right after chuckling at you for doing it.

OpenStudy (amistre64):

:) its like a cancer lol

OpenStudy (anonymous):

Almost, cancer can be cured, carelessness can't be.

OpenStudy (anonymous):

ok so can someone help me with the answer

OpenStudy (amistre64):

expand it out and slove for x :) 39^2 = x^2 +29^2 +2(29x) + x^2 the rest is just mathing it

OpenStudy (anonymous):

8.95